Q)The two blocks in an atwood machine have masses 2kg and 3 kg . find the work done by gravity during the fourth second after the system is released from rest?

Grade:12

1 Answers

Ramesh V
14 years ago

Let T be tension in rope

and a be accln of blocks

since 3kg is heavier one, it will tend to fall due to gravity

so, on balancing eqns, we have a= (m1-m2).g/(m1+m2)

    T = 2m1.m2.g/(m1+m2)

so T = 24 N and a = 2 m/sec2

at 4th second, after free fall , it falls a distance S = 1/2*2*16 = 16 m

work done by gravity = mgh = 3*10*16 = 480 J

----

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch